For the information given, find the values x,y and r. Clearly indicate the quadrant of the terminal side of theta, then state the values of the 6 trig functions of theta.
#61 sin (theta)= -(7/8) and secant (theta)< 0
I don't so much have trouble with figuring out the 6 functions, but rather how they determined in the answer that the quadrant of the terminal side of (theta) is in quadrant III.
Thank you in advance,
John

Answer by lwsshak3(11628)

Clearly indicate the quadrant of the terminal side of theta, then state the values of the 6 trig functions of theta. #61 sin (theta)= -(7/8) and secant (theta)< 0.
**
let x= theta
Given: sin x=-7/8 and sec x<0
This says sin x and sec x or sin x and cos x are in quadrants where both are negative. sin x and cos x are both negative only in quadrant III.
Trig functions as follows:
sin x=-7/8 (given)
cos x=-√15/8 (adjacent leg=√(8^2-7^2)=√64-49=√15)
tan x=7/√15=(7√15)/15
csc x=-8/7
sec x=-8√15=(-8√15)/15
cot x=√15/7